Write an equation for each parabola described below. Then draw the graph.
vertex (1,3)
directrix x =7/8
Could you show me the steps and process to get the answer?

Answer by stanbon(75887)

Write an equation for each parabola described below. Then draw the graph.
vertex (1,3)
directrix x =7/8
Could you show me the steps and process to get the answer?
-------
Note: The vertex is to the right of the directrix
so the parabola is opening to the right.
----
Form: (y-k)^2 = 4p(x-h)
----
p is the distance from the directrix to the vertex.
----
p = 1-(7/8) = 1/8
----
h = 1 and k = 3
-----------------------
Equation:
(y-3)^2 = 4(1/8)(x-1)
----
(y-3)^2 = (1/2)(x-1)
-------------------------
The graph is not a function so it cannot be shown on this site.
----
Sketch the vertical line x = 7/8
Plot the vertex at (1,3)
Sketch a parabola opening to the right.
===================
Cheers,
Stan H.
===================